# Parametricity via Cohesion

### v2

@@ 2.1 Path Types
In principle, given $a,b:A$, we could define the type of paths from $a$ to $b$ in $A$ as the type $\Sigma f:I\to A.(f~{}i_{0}=a)\times(f~{}i_{1}=b)$. However, experience with such a naïve formalization shows that it incurs a high cost of laborious transportations along identities that should be easy enough to infer automatically. Hence I instead follow the approach taken by Cubical Type Theory and related systems, and give an explicit axiomatization for path types, with corresponding rewrite rules to apply the associated equalities automatically:
@@
The introduction rule for path types corresponds to function abstraction.
@@
We may then postulate the usual $\beta$-law as an identity for this type, along with special identities for application to $i_{0}$ and $i_{1}$. All of these are made into rewrite rules, allowing Agda to apply them automatically, and thus obviating the need for excessive use of transport. [^5] We could additionally postulate an $\eta$-law for path types, analogous to the usual $\eta$-law for functions types; however, this is unnecessary for what follows, so I omit this assumption.
@@
With this formalization of path types in hand, we can straightforwardly formalize the equivalent formulation of weak connectedness of $I$ given above. For this purpose, we first define the map idToPath that takes an identification $a=b$ to a path from $a$ to $b$:
@@
A type $A$ is path-discrete if for all $a,b:A$, the map idToPath is an equivalence.
@@
which together imply that, if $A$ is discrete, then it is path-discrete.
@@
Although a full proof of canonicity is beyond the scope of this paper, I conjecture that adding these rules suffices to preserve canonicity.666perhaps not for Agda specifically, but for a more idealized system of type theory, such as Martin-Löf Type Theory I shall check a few specific cases of this conjecture later in the paper.
So much for the (weak) connectedness of $I$. Let us turn our attention now to the other property previously stipulated of $I$, namely strict bipointedness. As mentioned previously, we could simply postulate this stipulation directly as an axiom – however, for the purpose of proving parametricity theorems, a more prudent strategy is to instead formalize a class of $I$-indexed type families, whose computational behavior follows from this assumption (and which, in turn, implies it). Because these type families essentially correspond to the graphs of predicates and relations on arbitrary types, I refer to them as graph types.
@@ 2.2 Graph Types
The first projection from such a pair may be taken no matter what $i$ is, and reduces to the identity function when $i$ is $i_{0}$.
    g1fst :  {} {: Set ℓ} {: A  Set ℓ} (: I)             (: Gph1 i A B)  A    g1β1 :  {} {: Set ℓ} {: A  Set ℓ} (: I)               (: A)  (: (i ≡ i1)  B a)               g1fst i (g1pair {= B} i a b) ≡ a    {-# REWRITE g1β1 #-}    g1fst0 :  {} {: Set ℓ} {: A  Set ℓ}              (: Gph1 i0 A B)  g1fst {= B} i0 g ≡ g    {-# REWRITE g1fst0 #-}
@@
    g1snd :  {} {: Set ℓ} {: A  Set ℓ}             (: Gph1 i1 A B)  B (g1fst i1 g)    g1β2 :  {} {: Set ℓ} {: A  Set ℓ}            (: A)  (: (i1 ≡ i1)  B a)            g1snd (g1pair {= B} i1 a b) ≡ b refl    {-# REWRITE g1β2 #-}
It is straightforward to see that the inclusion of graph types makes strict bipointedness of $I$ provable, as follows:
strBpt : i0 ≡ i1  ⊥strBpt p = g1snd (transp  i  Gph1 i ⊤  _  ⊥)) p tt)
In fact, the converse holds under the assumption of univalence. Specifically, in the presence of univalence and the assumption of strict bipointedness for $I$, the type $\mathsf{Gph}^{1}~{}i~{}A~{}B$ may be regarded as a computationally convenient shorthand for the type $\Sigma x:A.(i=i1)\to Bx$, in much the same way as the type $\mathsf{Path}~{}A~{}a_{0}~{}a_{1}$ serves as shorthand for the type $\Sigma f:(\Pi i:I.Ai).(f~{}i_{0}=a_{0})\times(f~{}i_{1}=a_{1})$. This is due to the following equivalence:
@@
which is given by the map $\mathsf{fst}:(\Sigma x:A.(i_{0}=i_{1})\to Bx)\to A$ and which, under univalence, becomes an identity between its domain and codomain, thereby justifying the use of this and associated identities as axioms for graph types. I conjecture that moreover, taking these identities as rewrite rules is fully compatible with canonicity.777subject to the same caveat as before that this means canonicity for an idealized system of MLTT, and not Agda specifically
A few useful theorems concerning identities between elements of graph types are as follows, the latter of which is also made into a rewrite rule:
@@ 2.3 Parametricity via Sufficient Cohesion
I begin this section with an old chestnut of parametricity theorems – a proof that any polymorphic function of type (X : Set) → X → X must act as the identity function on any path-discrete type X.
@@
To prove the desired parametricity theorem for the type PolyId as above, we first prove a substitution lemma for this type, which can be stated as follows:
For any function α : PolyId and any path-discrete type A with a : A and a type family B : A → Set, there is a function B a → B (α A a).
The proof of this lemma is contained in the following submodule:
@@
The key step in the proof of this lemma is to construct a dependent path over the type family Gph1 i A B : I → Set as follows:
@@
And then since A is path-discrete, this path becomes an identity along which we can transport lemma1:
    substLemma : B (α A a)    substLemma = transp⁻¹ B (mkInv idToPath pdA lemma2) lemma1
From this substitution lemma, it straightforwardly follows that any element of PolyId must be extensionally equivalent to the identity function when evaluated at a path-discrete type.
@@
So far, we have proved only that the restrictions of elements of PolyId to path-discrete types are equivalent to that of the polymorphic identity function on path-discrete types. The theorem would then after all be trivial if it turned out that the only path-discrete types were (e.g.) those containing at most one element (i.e. the mere propositions, in the terminology of HoTT). To show that this is not the case, we make use of our assumption of (weak) connectedness for $I$, which we have already seen implies that every discrete type is path-discrete. One may think of the discrete types as forming their own distinct system of type theory which is then embedded in the broader system of cohesive types we have been working in, via the $\flat$ modality. We can then use the cohesive structure of the latter system to analyze properties of the embedded theory. Moreover, the subuniverse of discrete types contains many specific nontrivial types to which these theorems may be applied, e.g. – as in the following example – the type of Booleans Bool.
@@
Showing that Bool is discrete is a simple matter of pattern matching.
@@
This provides an opportunity to check some simple cases of the canonicity conjecture for the rewrite rules we have so far postulated. Specifically, we may check that the proof of polyId yields a canonical form (namely refl) when applied to canonical forms involving Booleans:
@@ 2.4 Parametricity & (Higher) Inductive Types
The foregoing proof of parametricity for the type of the polymorphic identity function remains ultimately a toy example. To demonstrate the true power of this approach to parametricity, I turn now to some more intricate examples of its use in proving universal properties for simplified representations of inductive types and higher inductive types.
In general, it is easy to write down what should be the recursion principle for an inductive type generated by a set of constructors, but harder (though feasible) to write down the corresponding induction principle. When one begins to consider more complex generalizations of inductive types, such as higher inductive types, inductive-inductive types, etc., these difficulties begin to multiply.
What would be ideal would be a way of deriving the induction principle for an inductive type from its recursor, hence requiring only the latter to be specified as part of the primitive data of the inductive type. However, in most systems of ordinary dependent type theory this is generally not possible [Geu01]. In HoTT, there is one way around this issue, due to Awodey, Frey & Speight [AFS18], whereby additional naturality constraints are imposed upon the would-be inductive type, that serve to make the corresponding induction principle derivable from the recursor. However, when one goes on to apply this technique to higher inductive types, which may specify constructs not only for elements of a type, but also for instances of its (higher) identity types, the complexity of these naturality conditions renders them impractical to work with. The ballooning complexity of these conditions is an instance of the infamous coherence problem in HoTT, whereby the complexity of coherence conditions for higher-categorical structures seemingly escapes any simple inductive definition.
As an alternative, let us consider ways of using parametricity to derive induction principles from recursion principles for inductive and higher inductive types, starting with the prototypical inductive type – the natural numbers $\mathbb{N}$.
First, we define the type of the recursor for $\mathbb{N}$.
@@
We may then postulate the usual constructors and identities for $\mathbb{N}$ [^8] Note that among the identities postulated as rewrite rules for $\mathbb{N}$ is its $\eta$-law, i.e. $\mathsf{rec}\mathbb{N}~{}n~{}\mathbb{N}~{}\mathsf{zero}~{}\mathsf{succ}=n$ for all $n:\mathbb{N}$. This identity will be important for the following derivation of the induction principle for $\mathbb{N}$, and postulating it as a rewrite rule, rather than an axiom, is thus necessary to preserve canonicity.
@@
From here, we may proceed as in the proof of parametricity for PolyId, by proving an analogous substitution lemma for Recℕ, following essentially the same steps:
@@
In order to apply this lemma to $\mathbb{N}$ itself, we must further postulate that $\mathbb{N}$ is path-discrete. [^9] One might think that we could show by induction that $\mathbb{N}$ is discrete, hence path-discrete by weak connectedness of $I$; however, since we have not yet proven induction for the above-given encoding of $\mathbb{N}$ internally, we must instead take path-discreteness of $\mathbb{N}$ as an additional axiom, from which induction on $\mathbb{N}$ will follow.
@@
Induction for $\mathbb{N}$ then follows as a straightforward consequence of the substitution lemma. [^10] Note that our use of the $\eta$-law for $\mathbb{N}$ as a rewrite rule is critical for the last step of this proof, since otherwise we would obtain not a proof of P n, but rather P (recℕ n ℕ succ zero).
@@
As in the case of the parametricity theorem for PolyId, we may test that the derived induction principle for $\mathbb{N}$ evaluates canonical forms to canonical forms. The following example tests this for the usual inductive proof that zero is an identity element for addition on the right (when addition is defined by induction on the left argument).
@@
Moving on from inductive types to higher inductive types, we may now consider deriving the induction principle for the circle $S^{1}$, defined to be the type freely generated by a single basepoint $\mathsf{base}:S^{1}$, with a nontrivial identification $\mathsf{loop}:\mathsf{base}=\mathsf{base}$. The recursor for $S^{1}$ thus has the following type:
@@
Then, just as before, we may postulate the corresponding constructors and $βη$-laws for $𝑆¹$.
@@
The proof of induction for $S^{1}$ is then in essentials the same as the one given above for $\mathbb{N}$. We begin by proving a substitution lemma for RecS¹, following exactly the same steps as in the proof of the corresponding lemma for Recℕ.
@@
We then postulate that $S^{1}$ is path-discrete, as before, in order to apply this lemma to $S^{1}$ itself.
@@
Then the desired induction principle for $𝑆¹$ follows straightforwardly.
@@ 3 Toward a Synthetic Theory of Parametricity
The theory so-far developed gives the rudiments of a synthetic framework for working in the internal language of a (weakly) sufficiently cohesive $\infty$-topos. This framework in turn proves capable of deriving significant parametricity results internally, with immediate applications in e.g. resolving coherence problems having to do with higher inductive types. It remains to be seen what further applications can be developed for this theory and its particular approach to parametricity. From this perspective, it is profitable to survey what other approaches there are to internalizing parametricity theorems in dependent type theory, and how they might be related to the one given in this paper.
@@ 3.1 Cohesion & Gluing
In recent years, there has been some related work toward a synthetic theory of parametricity in terms of the topos-theoretic construction of Artin Gluing. This approach, outlined initially by Sterling in his thesis [Ste22] and subsequently spearheaded by Sterling and his collaborators as part of the more general programme of Synthetic Tait Computability (STC), works in the internal language of a topos equipped with two (mere) propositions $L$ and $R$, that are mutually exclusive in that $L\wedge R\to\bot$. The central idea of this approach is to synthetically reconstruct the usual account of parametricity in terms of logical relations by careful use of the open and closed modalities induced by these propositions, which play a similar role in STC to the graph types introduced above. Using this approach, one can e.g. prove representation independence theorems for module signatures as in recent work by Sterling & Harper [SH21].
@@
On the other hand, there is no analogue in STC of the axiom of (weak) connectedness for the interval, and its consequent parametricity theorems, most notably the above-mentioned derivability of induction principles for higher inductive types. This suggests that in these latter results, the structure of sufficient cohesion and its relation to parametricity plays an essential role. Nonetheless, it remains interesting to consider how parametricity via cohesion and parametricity via gluing may yet prove to be related, and one may hope for a fruitful cross-pollination between these two theories.

### v3

@@ 2.1 Path Types
In principle, given $a,b:A$, we could define the type of paths from $a$ to $b$ in $A$ as the type $\Sigma f:I\to A.(f~{}i_{0}=a)\times(f~{}i_{1}=b)$. However, experience with such a naïve formalization shows that it incurs a high number of laborious transportations along equalities that should be easy enough to infer automatically. Hence I instead follow the approach taken by Cubical Type Theory and related systems, and give an explicit axiomatization for path types, with corresponding rewrite rules to apply the associated equalities automatically:
@@
The introduction rule for path types corresponds to function abstraction
@@
We may then postulate the usual $\beta$-law as an identity for this type, along with special identities for application to $i_{0}$ and $i_{1}$. All of these are made into rewrite rules, allowing Agda to apply them automatically, and thus obviating the need for excessive use of transport: [^3] We could additionally postulate an $\eta$-law for path types, analogous to the usual $\eta$-law for function types; however, this is unnecessary for what follows, and so I omit this assumption.
@@
With this formalization of path types in hand, we can straightforwardly formalize the equivalent formulation of weak connectedness of $I$ given above. For this purpose, we first define the map idToPath that takes an identification $a≡b$ to a path from $a$ to $b$:
@@
A type $A$ is path-discrete if for all $a,b:A$ the map idToPath is an equivalence:
@@
which together imply that, if $A$ is discrete, then it is path-discrete:
@@
Although a full proof of canonicity is beyond the scope of this paper, I conjecture that adding these rules suffices to preserve canonicity, and I verify a few concrete cases of this conjecture later in the paper.
So much for the (weak) connectedness of $I$; let us now turn our attention to the other property we had previously stipulated of $I$, namely its strict bipointedness. As mentioned previously, we could simply postulate this stipulation directly as an axiom – however, for the purpose of proving parametricity theorems, a more prudent strategy is to instead formalize a class of $I$-indexed type families, whose computational behavior follows from this assumption (and which, in turn, implies it). Because these type families essentially correspond to the graphs of predicates and relations on arbitrary types, I refer to them as graph types.
@@ 2.2 Graph Types
The first projection from such a pair may then be taken no matter what $i$ is, and reduces to the identity function when $i$ is $i_{0}$:
    g1fst :  {} {: Set ℓ} {: A  Set ℓ} (: I)             (: Gph1 i A B)  A    g1beta1 :  {} {: Set ℓ} {: A  Set ℓ} (: I)               (: A)  (: (i ≡ i1)  B a)               g1fst i (g1pair {= B} i a b) ≡ a    {-# REWRITE g1beta1 #-}    g1fst0 :  {} {: Set ℓ} {: A  Set ℓ}              (: Gph1 i0 A B)  g1fst {= B} i0 g ≡ g    {-# REWRITE g1fst0 #-}
@@
    g1snd :  {} {: Set ℓ} {: A  Set ℓ}             (: Gph1 i1 A B)  B (g1fst i1 g)    g1beta2 :  {} {: Set ℓ} {: A  Set ℓ}               (: A)  (: (i1 ≡ i1)  B a)               g1snd (g1pair {= B} i1 a b) ≡ b refl    {-# REWRITE g1beta2 #-}
It is straightforward to see that the inclusion of graph types makes strict bipointedness of the interval provable, as follows:
strBpt : (i0 ≡ i1)  ⊥strBpt p = g1snd (transp  i  Gph1 i ⊤  _  ⊥)) p tt)
And in fact, the converse holds under the assumption of univalence. Specifically, in the presence of univalence and the assumption of strict bipointedness for $I$, the type $\mathsf{Gph}^{1}~{}i~{}A~{}B$ may be regarded as a computationally convenient shorthand for the type $\Sigma x:A.(i=i_{1})\to Bx$, in much the same way as the type $\mathsf{Path}~{}A~{}a_{0}~{}a_{1}$ serves as shorthand for the type $\Sigma f:(\Pi i:I.Ai).(f~{}i_{0}=a_{0})\times(f~{}i_{1}=a_{1})$. This fact is due to the following equivalence
@@
which is given by the map $\mathsf{fst}:(\Sigma x:A.(i_{0}=i_{1})\to Bx)\to A$ and which, under univalence, becomes an identity between its domain and codomain, thereby justifying the use of this and associated identities as rewrite rules which, conjecturally, are fully compatible with canonictiy.
A few additional theorems, concerning identities between elements of graph types, are as follows, the latter of which I make into a rewrite rule for convenience:
@@ 2.3 Parametricity via Sufficient Cohesion
I begin this section with an old chestnut of parametricity theorems – a proof that any polymorphic function of type (X : Set) → X → X must be equivalent to the polymorphic identity function λ X → λ x → x.
@@
Before proceeding with this proof, however, it will be prudent to consider the meaning of this theorem in the context of the cohesive type theory we have so-far developed. Specifically, I wish to ask: over which types should the type variable X in the type (X : Set) → X → X be considered as ranging in the statement of this theorem? Although it is tempting to think that the answer to this question should be “all types” (or as close to this as one can get predicatively), if one considers the relation between our cohesive setup and Reynolds’ original setup of parametricity, a subtler picture emerges. A type, in our framework, corresponds not to a type in the object language of e.g. bare sets, but rather to an object of the cohesive topos used to interpret the parametric structure of this object language, e.g. the category of reflexive graphs. In this sense, we should expect the parametricity result for the type (X : Set) → X → X to generally hold only for those types corresponding to those in the object language, i.e. the discrete types. Indeed, the discrete types by construction are those which cannot distinguish elements of other types belonging to the same connected component, which intuitively corresponds to the fundamental idea of parametricity – that functions defined over these types must behave essentially the same for related inputs.
However, we cannot state this formulation of the theorem directly, since it would require us to bind $X$ as @♭ X : Set, which would kill all of the cohesive structure on Set and pose no restriction on the functions inhabiting this type. The solution, in this case, is to restrict the range of X to types which are path-discrete, since this requirement can be stated even for X not crisp.
To prove the desired parametricity theorem for the type PolyId as above, then, we first prove a substitution lemma as follows:
For any function α : PolyIdA and any path-discrete type A with a : A and a type family B : A → Set, there is a function B a → B(α A a)
@@
The key step in the proof of this lemma is to construct a “dependent path” over the type family Gph1 i A B : I → Set as follows:
@@
And then since A is path-discrete, this path becomes an equality along which we can transport lemma1:
    substLemma : B ( A a)
    substLemma = transp B (mkInv idToPath pdA lemma2) lemma1
From this substitution lemma, it straightforwardly follows that any element of PolyId must be extensionally equivalent to the polymorphic identity function when evaluated at a path-discrete type:
@@
Before we congratulate ourselves for proving this theorem, however, we ought to reflect on the significance of what we have proved. For we have proved only that the restrictions of elements of PolyId to path-discrete types are equivalent to that of the polymorphic identity function. The theorem would then after all be trivial if it turned out that the only path-discrete types were (e.g.) those containing at most one element (i.e. the mere propositions, in the terminology of HoTT). To show that this is not the case, we make use of our assumption of connectedness for $I$, which we have already seen implies that every discrete type is path-discrete. To give a concrete (non-trivial) instance of this, I now show that the type of Booleans is discrete (hence path-discrete) and use this to test the canonicity conjecture on a simple example:
@@
Showing that Bool is discrete is a simple matter of pattern matching
@@
We can use this to check that, in at least one specific case, the proof of polyId yields a canonical form (namely refl) when it is applied to canonical forms:
@@ 2.4 Parametricity & (Higher) Inductive Types
The foregoing proof of parametricity for the type of the polymorphic identity function remains ultimately a toy example. To demonstrate the true power of this approach to parametricity, I turn now to some more intricate examples of its use, in proving universal properties for simplified presentations of inductive types and higher inductive types.
In general, it is easy to write down what should be the recursion principle for an inductive type generated by some set of constructors, but harder (though feasible) to write down the corresponding induction principle. When one begins to consider more complex generalizations of inductive types, such as higher inductive types, inductive-inductive types, etc, these difficulties begin to multiply. What would be ideal would be a way of deriving the induction principle for an inductive type from its recursor, hence requiring only the latter to be specified as part of the primitive data of the inductive type. However, in most systems of ordinary dependent type theory this is generally not possible [7]. In HoTT, there is one way around this issue, due to Awodey, Frey & Speight [2], whereby additional naturality constraints are imposed upon the would-be inductive type, that serve to make the corresponding induction principle derivable from the recursor. However, when one goes on to apply this technique to higher inductive types, which may specify constructors not only for elements of a type, but also for instances of its (higher) identity types, the complexity of these naturality conditions renders them impractical to work with. The ballooning complexity of these conditions is an instance of the infamous coherence problem in HoTT, whereby the complexity of coherence conditions for higher-categorical structures seemingly escapes any simple inductive definition.
As an alternative, let us consider ways of using parametricity to derive induction principles for inductive and higher inductive types, starting with the prototypical inductive type, the natural numbers $\mathbb{N}$.
First, we define the type of the recursor for $\mathbb{N}$:
@@
We may then postulate the usual constructors and identities for $\mathbb{N}$ [^4] Note that among the identities postulated as rewrite rules for $\mathbb{N}$ is its $\eta$-law, i.e. $\mathsf{rec\mathbb{N}}~{}n~{}\mathbb{N}~{}\mathsf{zero}~{}\mathsf{succ}=n$ for all $n:\mathbb{N}$. This will be important for deriving the induction principle for $\mathbb{N}$.
@@
From here, we may proceed essentially as in the proof of parametricity for PolyId, by proving an analogous substitution lemma for Recℕ, following essentially the same steps:
@@
[…]
@@
Induction for $\mathbb{N}$ then follows as a straightforward consequence of the substitution lemma: [^5] Note that our use of the $\eta$-law for $\mathbb{N}$ as a rewrite rule is critical to the above proof, since otherwise in the last step, we would obtain not a proof of P n, but rather P (recℕ n ℕ succ zero).
@@
As in the case of the parametricity theorem for PolyId, we may test that the derived induction principle for $\mathbb{N}$ evaluates canonical forms to canonical forms. The following example tests this for the usual inductive proof that zero is an identity element for addition on the right:
@@
Moving on, then, from inductive types to higher inductive types, we may now consider deriving the induction principle for the circle $S^{1}$, defined to be the type freely generated by a single basepoint $\mathsf{base}:S^{1}$, with a nontrivial identification $\mathsf{loop}:\mathsf{base}=\mathsf{base}$. The recursor for $S^{1}$ thus has the following type:
@@
Then, just as before, we may postulate the corresponding constructors and $\beta\eta$-laws for $S^{1}$:
@@
The proof of induction for $S^{1}$ is then in essentials the same as the one given above for $\mathbb{N}$. We begin by proving a substitution lemma for RecS¹, following exactly the same steps as in the proof of the corresponding theorem for Recℕ:
@@
We then postulate that $S^{1}$ is path-discrete, as before, in order to apply this lemma to $S^{1}$ itself:
@@
And then the desired induction principle for $S^{1}$ follows straightforwardly:
@@ 3 Toward a synthetic theory of parametricity
The theory so-far developed essentially gives a synthetic framework for working in the internal language of a (weakly) sufficiently cohesive $\infty$-topos. This framework in turn proves capable of deriving significant parametricity results internally, with immediate applications in e.g. resolving coherence problems having to do with higher inductive types. It remains to be seen what further applications can be developed for this theory and its particular approach to parametricity. From this perspective, it is profitable to survey what other approaches there are to internalizing parametricity theorems in dependent type theory, and how they might be related to the one given in this paper.
@@ 3.1 Cohesion & Gluing
In recent years, there has been some related work toward a synthetic theory of parametricity in terms of the topos-theoretic construction of Artin Gluing. This approach, outlined initially by Sterling in his thesis [16] and subsequently spearheaded by Sterling and his collaborators as part of the more general programme of Synthetic Tait Computability (STC), works in the internal language of a topos equipped with two (mere) propositions $L$ and $R$, that are mutually exclusive in that $L\wedge R\to\bot$. The central idea of this approach is to synthetically reconstruct the usual account of parametricity in terms of logical relations by careful use of the open and closed modalities induced by these propositions, which play a similar role in STC to the graph types introduced above. Using this approach, one can e.g. prove representation independence theorems for module signatures as in recent work by Sterling & Harper [14].
@@
On the other hand, there is no analogue in STC of the axiom of connectedness for the interval, and its consequent parametricity theorems, most notably the above-mentioned derivability of induction principles for higher inductive types. This suggests that in these latter results, the structure of cohesion and its relation to parametricity plays an essential role. Nonetheless, it remains interesting to consider how parametricity via cohesion and parametricity via gluing may yet prove to be related, and one may hope for a fruitful cross-pollination between these two theories.